# LLaMA-Berry: Pairwise Optimization for O1-like Olympiad-Level Mathematical Reasoning

### v1

@@ D.1 Overview
Berry-Tree is an inference framework specifically designed for complex multi-model reasoning tasks, addressing the need to improve inference efficiency and accuracy in intricate tree search processes. This framework is particularly suited for large-scale reasoning tasks that involve the integration of multiple models and algorithms. By incorporating advanced tree search methods, robust concurrency handling, and lightweight inference engines, Berry-Tree significantly enhances inference speed and resource utilization. This report provides a detailed explanation of the system architecture and key technologies of Berry-Tree, along with preliminary performance evaluation results.
@@ D.2 System Architecture Overview
Figure 9 demonstrates the architecture of Berry-Tree which is divided into several layers, each handling different functional requirements. The Data Management Layer is responsible for the serialization and deserialization of data, ensuring efficient data read/write operations across models and systems. The Tree Search Methods Layer incorporates MCTS (Monte Carlo Tree Search), ToT (Tree of Thoughts), and A* algorithms to optimize the inference process and explore multiple reasoning paths. Additionally, Berry-Tree includes a Reliability Layer, which ensures load balancing and failover support in highly concurrent scenarios, guaranteeing the stability of inference tasks. Finally, the Inference Engines Layer integrates efficient inference engines such as vLLM, FastAPI, and HuggingFace TGI to enable parallelized and efficient task processing.
@@ D.3 Key Technical Components
Data Management: Berry-Tree employs serialization and deserialization techniques, specifically using formats like JSON and CSV, to efficiently store and transfer checkpoint data from tree search reasoning processes. The framework stores this data along with hash values to ensure integrity and allows for quick restoration of tree search states in memory when needed. Furthermore, Berry-Tree leverages the HuggingFace Datasets library to handle core data inputs for both training and inference. It supports seamless loading of datasets such as GSM8K and MATH from the HuggingFace Hub, enhancing its flexibility and ensuring compatibility with diverse reasoning tasks.
Tree Search Methods: Berry-Tree support multiple tree search algorithms, with MCTS being central to handling large-scale complex reasoning tasks by leveraging random simulation and statistical analysis to optimize the search space. The Tree of Thoughts (ToT) further extends the exploration depth and breadth of reasoning paths, helping the system manage uncertainty. A* is applied in pathfinding tasks, offering efficient heuristic search capabilities.
Reliability Design: To ensure the stability and continuity of inference tasks, Berry-Tree incorporates load balancing and failover mechanisms. During high-concurrency operations, the load balancer distributes workloads across different servers, preventing server overload. The failover mechanism ensures that tasks can seamlessly transition to backup servers in case of partial server failures.
Server Architecture: The framework’s server architecture is divided into two segments: one dedicated to executing Large Language Model (LLM) inference and the other designated explicitly for handling PPRM (Pairwise Preference Reward Model) and EBC (Enhanced Borda Count) processes. This modular design allows the framework to allocate computational resources flexibly, improving overall efficiency.
Inference Engines: Inference engines of Berry-Tree include VLLM (a lightweight language model inference engine), FastAPI (a high-performance web framework), and HuggingFace TGI (Text Generation Inference engine). These engines collectively enable the system to maintain high efficiency and stability while handling multi-model inference tasks, with robust support for high-concurrency demands.
@@ D.4 Preliminary Performance Evaluation
We conduct 16 MCTS rollouts to parallelize the inference tasks via Berry-Tree. The results indicate that the total inference time is 1 hour and 25 minutes, with an average inference time of approximately 3.87 seconds per sample. These results demonstrate strong parallel inference capabilities of Berry-Tree under the given hardware configuration.
This preliminary performance evaluation highlights the efficiency of Berry-Tree, particularly in handling complex reasoning tasks. Although these results are based on initial experiments, future evaluations with varying hardware configurations and more complex task scenarios could provide deeper insights into the system’s performance.
D.5 System Scalability and Optimization
The modular design of Berry-Tree allows for future scalability and optimization. For instance, the system can be flexibly integrated with additional inference engines, support alternative tree search methods, or accommodate more diverse language models. This flexibility makes it adaptable to various applications, such as natural language reasoning, mathematical problem-solving, and pathfinding tasks. Future optimization directions will include enhancing inference speed, improving concurrency handling, and strengthening fault-tolerance and load management mechanisms to ensure stability and reliability in large-scale deployment scenarios.
Appendix E Convergence Analysis of the Enhanced Borda Count (EBC) Method
@@
In this appendix section, we provide a rigorous mathematical proof that the quantile scores evaluated by the Enhanced Borda Count (EBC) method converge to the true quantile scores of solutions in the actual quality distribution as the number of samples increases.
E.1 Definitions and Assumptions
Solution Set: Let $A=\{a_{1},a_{2},\ldots,a_{n}\}$ be a finite set of $n$ solutions (answers).
@@ E.1 Definitions and Assumptions
True Quality Scores: Each solution $a_{i}$ has a true quality score $Q^{*}(a_{i})\in\mathbb{R}$, drawn from a continuous distribution $P_{Q}$.
@@
True Ranking: The true ranking $R^{*}$ is induced by ordering the solutions in decreasing order of their true quality scores $Q^{*}(a_{i})$.
@@
True Pairwise Preference Probabilities: The true probability that solution $a_{i}$ is preferred over $a_{j}$ is defined as:
$P^{*}(a_{i}\succ a_{j})=\mathbb{P}(Q^{*}(a_{i})>Q^{*}(a_{j}))+\tfrac{1}{2}\mathbb{P}(Q^{*}(a_{i})=Q^{*}(a_{j})).$ (11)
@@
$P^{*}(a_{i}\succ a_{j})=\begin{cases}1,&\text{if }Q^{*}(a_{i})>Q^{*}(a_{j}),\\ 0,&\text{if }Q^{*}(a_{i})<Q^{*}(a_{j}).\end{cases}$ (12)
Estimated Preference Probabilities: For $T$ independent samples, the estimated preference probability is:
$P_{T}(a_{i}\succ a_{j})=\frac{1}{T}\sum_{t=1}^{T}X_{t}^{(i,j)},$ (13)
@@
Convergence Assumption: As $T\to\infty$, $P_{T}(a_{i}\succ a_{j})$ converges almost surely to $P^{*}(a_{i}\succ a_{j})$.
Preference Matrix: Construct the estimated preference matrix $M_{T}$ and the true preference matrix $M^{*}$ as:
$M_{T}[i,j]=\begin{cases}1,&\text{if }P_{T}(a_{i}\succ a_{j})\geq 0.5,\\ 0,&\text{otherwise.}\end{cases}\quad M^{*}[i,j]=\begin{cases}1,&\text{if }P^{*}(a_{i}\succ a_{j})=1,\\ 0,&\text{if }P^{*}(a_{i}\succ a_{j})=0.\end{cases}$ (14)
Transitive Closure: Let $C_{T}$ and $C^{*}$ be the transitive closures of $M_{T}$ and $M^{*}$, respectively.
Borda Counts: Compute the Borda count for each solution:
$B_{T}(a_{i})=\sum_{j\neq i}C_{T}[i,j],\quad B^{*}(a_{i})=\sum_{j\neq i}C^{*}[i,j].$ (15)
Estimated Ranking and Quantile Scores:
@@
    $Q_{g}(a_{i})=1-\frac{\text{rank}_{T}(a_{i})-1}{n-1}.$ (16)
@@
    $Q^{*}_{g}(a_{i})=1-\frac{\text{rank}^{*}(a_{i})-1}{n-1}.$ (17)
E.2 Objective
@@
$\lim_{T\to\infty}Q_{g}(a_{i})=Q^{*}_{g}(a_{i}),\quad\forall a_{i}\in A.$ (18)
E.3 Proof
Step 1: Convergence of Estimated Preference Probabilities
@@ E.3 Proof
$\lim_{T\to\infty}P_{T}(a_{i}\succ a_{j})=P^{*}(a_{i}\succ a_{j})\quad\text{almost surely}.$ (19)
Step 2: Convergence of Preference Matrix $M_{T}$ to $M^{*}$
@@
Step 3: Convergence of Transitive Closure $C_{T}$ to $C^{*}$
@@
$C_{T}\to C^{*}\quad\text{as }T\to\infty,$ (21)
@@
Step 4: Convergence of Borda Counts $B_{T}(a_{i})$ to $B^{*}(a_{i})$
@@
Step 5: Convergence of Estimated Ranking $R_{T}$ to True Ranking $R^{*}$
@@
$R_{T}\to R^{*}\quad\text{as }T\to\infty,$ (23)
@@
Step 6: Convergence of Quantile Scores $Q_{g}(a_{i})$ to $Q^{*}_{g}(a_{i})$
@@
$Q_{g}(a_{i})=1-\frac{\text{rank}_{T}(a_{i})-1}{n-1}\to Q^{*}_{g}(a_{i})=1-\frac{\text{rank}^{*}(a_{i})-1}{n-1},$ (24)
@@ Conclusion
$\lim_{T\to\infty}Q_{g}(a_{i})=Q^{*}_{g}(a_{i}),\quad\forall a_{i}\in A,$ (25)
@@
E.4 Finite Sample Analysis and Convergence Rate
Lemma 1: Hoeffding’s Inequality for Preference Probability Estimates
@@ E.4 Finite Sample Analysis and Convergence Rate
$\mathbb{P}\left(\left|P_{T}(a_{i}\succ a_{j})-P^{*}(a_{i}\succ a_{j})\right|\geq\epsilon\right)\leq 2\exp(-2T\epsilon^{2}).$ (26)
Lemma 2: Uniform Convergence over All Pairs
@@
$\mathbb{P}\left(\exists(i,j):\left|P_{T}(a_{i}\succ a_{j})-P^{*}(a_{i}\succ a_{j})\right|\geq\epsilon\right)\leq N\cdot 2\exp(-2T\epsilon^{2}).$ (27)
@@
$T\geq\frac{1}{2\epsilon^{2}}\ln\left(\frac{2N}{\delta}\right).$ (28)
Lemma 3: Correctness of Preference Matrix with High Probability
@@
$\left|P_{T}(a_{i}\succ a_{j})-P^{*}(a_{i}\succ a_{j})\right|<\epsilon,$ (29)
@@
Lemma 4: Probability of Correct Ranking
@@
Convergence Rate Analysis
@@
$T\geq\frac{1}{2\epsilon^{2}}\ln\left(\frac{n(n-1)}{\delta}\right).$ (30)
@@
E.5 Addressing Practical Considerations
@@
    $\left|P^{*}(a_{i}\succ a_{j})-0.5\right|\geq m.$ (32)
@@
    $\mathbb{P}\left(M_{T}[i,j]\neq M^{*}[i,j]\right)\leq 2\exp(-2Tm^{2}).$ (33)
@@
    $T\geq\frac{1}{2m^{2}}\ln\left(\frac{n(n-1)}{\delta}\right).$ (34)
@@
E.6 Final Conclusion
We have provided a rigorous proof that the EBC method’s estimated quantile scores $Q_{g}(a_{i})$ converge almost surely to the true quantile scores $Q^{*}_{g}(a_{i})$ as the number of samples $T$ approaches infinity. The finite sample analysis demonstrates that the convergence rate depends logarithmically on the number of solutions and inversely on the square of the margin $m$ between preference probabilities.
Appendix F Pseudo-code of main Algorithms
Data: $M$ (Binary Reward Matrix),$P$ (Pairwised Preference Reward model)

### v2

@@ D.1 Overview
Berry-Tree is an inference framework specifically designed for complex multi-model reasoning tasks, addressing the need to improve inference efficiency and accuracy in intricate tree search processes of LLM’s mathematical reasoning. This framework is particularly suited for large-scale reasoning tasks that involve the integration of multiple models and algorithms. By incorporating advanced tree search methods especially Monte Carlo Tree Search (MCTS), robust concurrency handling, and high-performance inference engines, Berry-Tree significantly enhances inference speed and resource utilization of LLM’s mathematical reasoning process. This section provides a explanation of the system architecture and key technologies of Berry-Tree, along with a preliminary performance evaluation results.
@@ D.2 System Architecture Overview
Figure A3 demonstrates the architecture of Berry-Tree which is divided into several layers, each handling different functional requirements. The Data Management Layer is responsible for the serialization and deserialization of data, ensuring efficient data read/write operations across models and systems and the ablity of recovering the search process from serialized data. The Tree Search Methods Layer incorporates MCTS (Monte Carlo Tree Search), ToT (Tree of Thoughts), and A* algorithms to optimize the inference process and explore multiple reasoning paths. Additionally, Berry-Tree includes a Reliability Layer, which ensures load balancing and failover support in highly concurrent scenarios, guaranteeing the stability of inference tasks. Finally, the Inference Engines Layer integrates efficient inference engines such as vLLM, FastAPI, and HuggingFace TGI to enable parallelized and efficient task processing.
@@ D.3 Key Technical Components
Figure A3: Architecture design of Berry-Tree
Data Management. Berry-Tree employs serialization and deserialization techniques, specifically using formats including JSON and CSV, to efficiently store, transfer and recover checkpoint data from tree search reasoning processes. The framework stores this data along with hash values to ensure integrity and allows for quick restoration of tree search states in memory when needed. Furthermore, Berry-Tree leverages the HuggingFace Datasets library to handle core dataset inputs for both training and inference. It supports seamless loading of benchmark datasets such as GSM8K and MATH from the HuggingFace Hub, enhancing its flexibility and ensuring compatibility with diverse reasoning tasks.
Tree Search Methods. Berry-Tree support multiple tree search algorithms, with MCTS being central to handling large-scale complex reasoning tasks by leveraging random simulation and statistical analysis to optimize the search space. The Tree of Thoughts (ToT) extends the exploration depth and breadth of reasoning paths, helping the system manage uncertainty. And A* can offer a efficient heuristic search capabilities.
Reliability Design. To ensure the stability and continuity of inference tasks, Berry-Tree incorporates load balancing and failover mechanisms. During high-concurrency operations, the load balance componet distributes workloads across different inference servers, preventing server overload. The failover mechanism ensures that tasks can seamlessly recover and transition to backup servers in case of partial server failures.
Server Architecture. The framework’s server architecture is divided into two segments: one dedicated to executing Large Language Model (LLM) inference and the other designated explicitly for handling PPRM (Pairwise Preference Reward Model) with EBC (Enhanced Borda Count) method. This modular design allows the framework to allocate computational resources flexibly, improving overall efficiency.
Inference Engines. Inference engines of Berry-Tree include VLLM , HuggingFace Transformers warpped by FastAPI, and HuggingFace TGI. These engines collectively enable the system to maintain high efficiency and stability while handling multi-model inference tasks, with robust support for high-concurrency demands.
@@ D.4 Preliminary Performance Evaluation
We conduct 16 rollouts to parallelize the inference tasks of LLaMA-Berry via Berry-Tree. The results indicate that the total inference time is 1 hour and 25 minutes, with an average inference time of approximately 3.87 seconds per sample. These results demonstrate a strong parallel inference capabilities of Berry-Tree under the given hardware configuration.
Appendix E Scaling Study on Inference-time Token Overheads
Figure A4: Average token consumption comparison across datasets, error bar stands for standard deviation.
Figure A5: Average token consumption for LLaMA-3.1-8B-Instruct across olympiad-level datasets, error bar stands for standard deviation.
Comparing LLaMA-3-8B-Instruct and LLaMA-3.1-8B-Instruct across GSM8K, GSMHARD, and MATH500, LLaMA-3.1-8B-Instruct consistently consumes more tokens across all categories—Solutions, Critiques, and Overall. This Figure A4 result reflects LLaMA-3.1-8B-Instruct’s tendency to generate more detailed and comprehensive outputs with increased overhead. While this likely improves solution quality, it also introduces greater resource demands and variability in token usage, highlighting a trade-off between accuracy and computational resources. LLaMA-3.1-8B-Instruct is thus better suited for tasks prioritizing precision over speed. As shown in Figure  A5, Token overheads during inference also scale with task difficulty across different olympiad-level benchmarks with LLaMA-3.1-8B-Instruct. AIME2024 exhibits the highest token consumption with significant variability, reflecting the complexity of its solution paths. In contrast, GPQA Diamond shows lower overall overhead, while AMC2023 falls in between, with moderate token consumption and less variability than AIME2024 but still notable.
Appendix F Future Work
LLaMA-Berry holds significant potential for further development. First, we plan to enhance its multimodal capabilities, enabling it to better handle complex problems like VQA and mathematic geometric problems that require visual, auditory, or even tactile perception. For instance, in tasks involving visual reasoning, such as geometric problems, LLaMA-Berry could be extended to integrate image recognition and analysis capabilities, thereby assisting in solving challenges related to spatial relationships and shape recognition. Furthermore, we aim to generalize its application across other scientific domains, improving its performance in disciplines such as physics, chemistry, and biology. In physics, for example, it could be utilized to address complex micro- and macroscopic dynamics problems; in chemistry, it may assist in molecular structure prediction and drug design; and in biology, it could potentially be used for genome analysis and disease prediction. Moreover, LLaMA-Berry can also offer technical support for other interdisciplinary fields, such as meteorology, environmental science, and materials science, by integrating multimodal data to enhance predictive accuracy. At the same time, we will focus on how LLaMA-Berry can further enhance AI safety. For instance, LLaMA-Berry could be leveraged to design more robust safety and risk assessment mechanisms. Besides, the generation of responses could be guided by a safety-trained PPRM to produce outputs with higher safety standards. These advancements not only pave the way for broader scientific applications of LLaMA-Berry but also offer new possibilities for enhancing AI safety and promoting its widespread adoption.
Appendix G Case Study
The prompts utilized in LLaMA-Berry for the LLaMA-3.1-8B-Instruct model are presented in Figure A6. Additionally, Figure A7 provides a detailed breakdown of problem-solving examples derived from the GSM8K dataset.
Figure A6: Prompts for LLaMA-3.1-8B-Instruct
Figure A7: Problem-solving example
Appendix H Convergence Analysis of the Enhanced Borda Count (EBC) Method
@@
In this appendix, we present a formal discussion on how the quantile scores, as evaluated by the Enhanced Borda Count (EBC) method, converge to the true quantile scores of solutions within the actual quality distribution as the number of samples increases.
H.1 Definitions and Assumptions
Solution Set. Let $A=\{a_{1},a_{2},\ldots,a_{n}\}$ be a finite set of $n$ solutions (answers).
@@ H.1 Definitions and Assumptions
True Quantile Scores. Each solution $a_{i}$ has a true quality score $Q^{*}(a_{i})\in\mathbb{R}$, drawn from a continuous distribution $P_{Q}$.
@@
True Ranking. The true ranking $R^{*}$ is induced by ordering the solutions in decreasing order of their true quantile scores $Q^{*}(a_{i})$.
@@
True Pairwise Preference Probabilities. The true probability that solution $a_{i}$ is preferred over $a_{j}$ is defined as:
$P^{*}(a_{i}\succ a_{j})=$ (A1)
$\mathbb{P}(Q^{*}(a_{i})>Q^{*}(a_{j}))+\tfrac{1}{2}\mathbb{P}(Q^{*}(a_{i})=Q^{*}(a_{j})).$
@@
$P^{*}(a_{i}\succ a_{j})=\begin{cases}1,&\text{if }Q^{*}(a_{i})>Q^{*}(a_{j}),\\ 0,&\text{if }Q^{*}(a_{i})<Q^{*}(a_{j}).\end{cases}$ (A2)
Estimated Preference Probabilities. For $T$ independent samples, the estimated preference probability is:
$P_{T}(a_{i}\succ a_{j})=\frac{1}{T}\sum_{t=1}^{T}X_{t}^{(i,j)},$ (A3)
@@
Convergence Assumption. As $T\to\infty$, $P_{T}(a_{i}\succ a_{j})$ converges almost surely to $P^{*}(a_{i}\succ a_{j})$.
Preference Matrix. Construct the estimated preference matrix $M_{T}$ and the true preference matrix $M^{*}$ as:
$M_{T}[i,j]=\begin{cases}1,&\text{if }P_{T}(a_{i}\succ a_{j})\geq 0.5,\\ 0,&\text{otherwise.}\end{cases}$ (A4)
$M^{*}[i,j]=\begin{cases}1,&\text{if }P^{*}(a_{i}\succ a_{j})=1,\\ 0,&\text{if }P^{*}(a_{i}\succ a_{j})=0.\end{cases}$
Transitive Closure. Let $C_{T}$ and $C^{*}$ be the transitive closures of $M_{T}$ and $M^{*}$, respectively.
Borda Counts. Compute the Borda count for each solution:
$B_{T}(a_{i})=\sum_{j\neq i}C_{T}[i,j],\quad B^{*}(a_{i})=\sum_{j\neq i}C^{*}[i,j].$ (A5)
Estimated Ranking and Quantile Scores.
@@
    $Q_{g}(a_{i})=1-\frac{\text{rank}_{T}(a_{i})-1}{n-1}.$ (A6)
@@
    $Q^{*}_{g}(a_{i})=1-\frac{\text{rank}^{*}(a_{i})-1}{n-1}.$ (A7)
H.2 Objective
@@
$\lim_{T\to\infty}Q_{g}(a_{i})=Q^{*}_{g}(a_{i}),\quad\forall a_{i}\in A.$ (A8)
H.3 Proof
Convergence of Estimated Preference Probabilities.
@@ H.3 Proof
$\lim_{T\to\infty}P_{T}(a_{i}\succ a_{j})=P^{*}(a_{i}\succ a_{j})\quad\text{almost surely}.$ (A9)
Convergence of Preference Matrix $M_{T}$ to $M^{*}$.
@@
Convergence of Transitive Closure $C_{T}$ to $C^{*}$.
@@
$C_{T}\to C^{*}\quad\text{as }T\to\infty,$ (A11)
@@
Convergence of Borda Counts $B_{T}(a_{i})$ to $B^{*}(a_{i})$.
@@
Convergence of Estimated Ranking $R_{T}$ to True Ranking $R^{*}$.
@@
$R_{T}\to R^{*}\quad\text{as }T\to\infty,$ (A13)
@@
Convergence of Quantile Scores $Q_{g}(a_{i})$ to $Q^{*}_{g}(a_{i})$.
@@
$Q_{g}(a_{i})=1-\frac{\text{rank}_{T}(a_{i})-1}{n-1}\to$ (A14)
$Q^{*}_{g}(a_{i})=1-\frac{\text{rank}^{*}(a_{i})-1}{n-1}$
@@ Conclusion.
$\lim_{T\to\infty}Q_{g}(a_{i})=Q^{*}_{g}(a_{i}),\quad\forall a_{i}\in A,$ (A15)
@@
H.4 Finite Sample Analysis and Convergence Rate
Lemma 1: Hoeffding’s Inequality for Preference Probability Estimates.
@@ H.4 Finite Sample Analysis and Convergence Rate
$\mathbb{P}\left(\left|P_{T}(a_{i}\succ a_{j})-P^{*}(a_{i}\succ a_{j})\right|\geq\epsilon\right)$ (A16)
$\leq 2\exp(-2T\epsilon^{2}).$
Lemma 2: Uniform Convergence over All Pairs.
@@
$\mathbb{P}\left(\exists(i,j):\left|P_{T}(a_{i}\succ a_{j})-P^{*}(a_{i}\succ a_{j})\right|\geq\epsilon\right)$ (A17)
$\leq N\cdot 2\exp(-2T\epsilon^{2}).$
@@
$T\geq\frac{1}{2\epsilon^{2}}\ln\left(\frac{2N}{\delta}\right).$ (A18)
Lemma 3: Correctness of Preference Matrix with High Probability.
@@
$\left|P_{T}(a_{i}\succ a_{j})-P^{*}(a_{i}\succ a_{j})\right|<\epsilon,$ (A19)
@@
Lemma 4: Probability of Correct Ranking.
@@
Convergence Rate Analysis.
@@
$T\geq\frac{1}{2\epsilon^{2}}\ln\left(\frac{n(n-1)}{\delta}\right).$ (A20)
@@
H.5 Addressing Practical Considerations
@@
    $\left|P^{*}(a_{i}\succ a_{j})-0.5\right|\geq m.$ (A22)
@@
    $\mathbb{P}\left(M_{T}[i,j]\neq M^{*}[i,j]\right)\leq 2\exp(-2Tm^{2}).$ (A23)
@@
    $T\geq\frac{1}{2m^{2}}\ln\left(\frac{n(n-1)}{\delta}\right).$ (A24)
@@
H.6 Final Conclusion
We have provided a formal discussion showing that the estimated quantile scores $Q_{g}(a_{i})$, obtained through the EBC method, converge almost surely to the true quantile scores $Q^{*}_{g}(a_{i})$ as the number of samples $T$ approaches infinity. The finite sample analysis demonstrates that the convergence rate depends logarithmically on the number of solutions and is inversely proportional to the square of the margin $m$ between preference probabilities.
Appendix I Pseudo-code of main Algorithms
Output: Ranked solution list $S$
1
2Initialize search tree $\mathcal{T}$ with root node $s_{0}$;
3 while number of nodes $N(\mathcal{T})<N_{max}$ do
4       ———Selection Phase———
6      ———Expansion Phase———
7       Expand $s_{i}$ by generating a successor node $s^{\prime}$ using the rewriting process $R(s_{i},c_{i})$, where $c_{i}=C(s_{i})$ is a critique of the current state;
8       Add the new node $s^{\prime}$ to $\mathcal{T}$;
9
10      ———Evaluation Phase———
11
$Q(s^{\prime})=\alpha Q_{g}(s^{\prime})+(1-\alpha)Q_{l}(s^{\prime})$
where $Q_{g}(s^{\prime})$ is the global value from the win-loss matrix $M$ and $Q_{l}(s^{\prime})$ is the local value from adjacent nodes in $\mathcal{T}$;
13
14      ———Backpropagation Phase———
15
$Q(s_{i})=(1-\gamma)Q(s_{i})+\gamma Q(s^{\prime})$
17      ———Check for tree growth limit———
18       if $N(\mathcal{T})\geq N_{max}$ then
19             break;
20
21       end if
22
23 end while
return Ranked solution list $S$
Data: $M$ (Binary Reward Matrix),$P$ (Pairwise Preference Reward model)